ok my question is, plot atleast five points for each equations
x+4y = 5
I have plugged in every possible number and I cant seem to get it

Answer by troyapplehelen(46)

sorry my mistake x+4y=5 is 4y=5-x, y=(5/4)-(x/4)
now, this a line with slope of -1/4, x-intercept is (5,0), y-intercept (0,5/4)
draw line passing these points in a graph paper.
to find other points, substitute x=1, then y=(5/4)-(1/4)= 4/4 =1 so (1,1)
x=3, then y=(5/4)-(3/4)= 2/4 = 1/2 so, (3,1/2)
x=-1, then y=(5/4)-(-1/4) = (5/4)+(1/4)
= 6/4 (=3/2)
so, (-1,3/2).